Circular No. 6128

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY NOVA IN SCORPIUS B. A. Harmon, S. N. Zhang, G. J. Fishman, C. A. Wilson, W. S. Pa-

ciesas, and B. C. Rubin, report for the Compton Gamma Ray Observatory BATSE Team: “This object (GRO J1655–40) has become detectable again, varying between 150–250 mCrab in the band 20–100 keV since Jan. 12. The previous hard x-ray outburst (IAUC 6101, 6106) ended in late December 1994. Emission is seen to at least 200 keV.”

SATURN J. Lecacheux, Observatoire de Meudon; and A. Sanchez-Lavega, Bilbao

University, report: “From CCD imaging of Saturn on 1994 Nov. 22.8 and Dec. 16.8–17.8 UT with the Pic du Midi 1-m telescope, we derive a drift rate of +9^{o}.9/day in longitude (system I) for the bright yellowish equatorial disturbance observed by the Hubble Space Telescope on Dec. 1 (confirming our previous report, IAUC 6079). We have also recorded another major bright feature, 135^{o} eastward, with similar latitude and motion; the wind velocity appears slower by 100 m/s than normally expected at this latitude.”

CATALOGUE OF COMETARY ORBITS 1995 A new edition of the Catalogue of Cometary Orbits, complete through

the end of 1994, is now available. It utilizes the new system of cometary designations and includes detailed cross-references between this system and the old system (both year/letter and year/Roman numeral). Other changes from the previous editions are that the general catalogue lists the singleapparition and multiple-apparition (i.e., numbered) comets separately (the latter chronologically by comet number), and that all the information (including the literature reference) about a particular orbit is placed on a single line spanning two facing pages. Separate entries are given for individual components of several split comets, yielding a total of 1472 entries for 1444 cometary apparitions referring to 878 different comets, 184 of them of short period. As an entirely new feature, there is a special tabulation giving osculating elements for the 116 numbered comets (except for five that are lost) for the epochs 1995 Mar. 24 and Oct. 10. The 108-page Catalogue sells for $20.00 ($30.00 for airmail delivery outside North America). There are also diskette and e-mail editions with the basic orbital information, and these cost $110.00 ($120.00 by airmail) and $50.00, respectively.

1995 January 27 Daniel W. E. Green